3. Below you are provided with the sum of squares, the variance, or the standard deviation for a sample, as well as the samplPlease help me with all 4.
Answer #1

s​​​​​​2 = SS/(n-1)

A) s​​​​​​2 = 2500/25 = 100

s= √100 = 10

B) s​​​​​​2 =25

SS= 4 X 25 = 100

C) SS= 100/50 = 2

s= √100 = 10

D) s​​​​​​2 =0.25

SS= 0.25/120= 0.00208

please like ??